Ten pounds of mixed birdseed sells for $7.30 per pound. The mixture is obtained from two kinds of birdseed, with one variety priced at $5.74 per pound and the other at $8.86 per pound. How many pounds of each variety of birdseed are used in the mixture?
how much pounds at $5.74 per pound?
how much pounds at $8.86 per pound?


Answer by mananth(16946)

--------$ per kg----------------quantity
Bird seed I 5.74 ----------------x pounds
Bird seed II 8.86 ----------------10-x pounds
Total 7.30% ---------------- 10 \
5.74x+8.86(10-x)=7.3*10
5.74x+88.6-8.86 x=73
5.74x-8.86x=73-88.6
-3.12x =-15.6
/-3.12
x=5 pounds of 5.74%
5 pounds of 8.86%